Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,000 helmets, using 2,010 kilograms of plastic. The plastic cost the company $17,286.

According to the standard cost card, each helmet should require 0.58 kilograms of plastic, at a cost of $9.00 per kilogram.

Required:

1. What is the standard quantity of kilograms of plastic (SQ) that is allowed to make 3,000 helmets?

2. What is the standard materials cost allowed (SQ × SP) to make 3,000 helmets?

3. What is the materials spending variance?

4. What is the materials price variance and the materials quantity variance?
Answer #1

Standard Quantity allowed

= 3,000 * 0.58

= 1740

Standard materials cost

= 1740 * 9

= 15,660

Materials spending variance = 15,660 - 17,286

= 1626 U

Materials Price Variance = (SP-AP) * AQ

= (9 - 17286/2010) * 2010

= 804 F

Materials Quantity Variance = (SQ-AQ) * SP

= (1740 - 2010) * 9

= 2430 U